My question is;
Michael and Eduardo are selling boy scouts cookies (a recent invention to try and compete with girl scouts). the boys sold 57 boxes of cookies, and earned $195. If buttery tea dippers costs $2.00 per box and Nutty cost $3.50 per box, how many boxes of each did they sell?

Answer by checkley79(3341)

X+Y=57
X=57-Y
3.5X+2Y=195
3.5(57-Y)+2Y=195
199.5-3.5Y+2Y=195
-1.5Y=195-199.5
-1.5Y=-4.5
Y=-4.5/-1.5
Y=3 BOXES OF $200 COOKIES WERE SOLD.
X+3=57
X=57-3
X=54 BOXES OF $3.50 COOKIES WERE SOLD.
PROOF:
3.5*54+2*3=195
189+6=195
195=195